OpenStudy (anonymous):

How many ways can you purchase 5 CDs if there are 6 to choose from, 2 cassettes if there are 5 to choose from, and 4 DVDs if there are 8 to choose from?

OpenStudy (anonymous):

altogether?

OpenStudy (anonymous):

\[\dbinom{6}{5}\times \dbinom{5}{2}\times \dbinom{8}{4}\] is how you do this

OpenStudy (anonymous):

\[\dbinom{6}{5}=6\] \[\dbinom{5}{2}=\frac{5\times 4}{2}=10\] \[\dbinom{8}{4}=\frac{8\times 7\times 6\times 5}{4\times 3\times 2}=2\times 7\times 5=70\]